# REP: Resource-Efficient Prompting for Rehearsal-Free Continual Learning

### v2

@@ Abstract
Recent rehearsal-free methods, guided by prompts, generally excel in vision-related continual learning (CL) scenarios with continuously drifting data. To be deployable on real-world devices, these methods must maintain high resource efficiency during training. In this paper, we introduce Resource-Efficient Prompting (REP), which targets improving the resource efficiency of prompt-based rehearsal-free methods. Our key focus is on avoiding catastrophic trade-offs with accuracy while trimming computational and memory costs during prompt learning. We achieve this by exploiting swift prompt selection that enhances input data using a carefully provisioned model, and by developing adaptive token merging (AToM) and layer dropping (ALD) algorithms for the prompt updating stage. AToM and ALD perform selective skipping across the data and model-layer dimensions without compromising task-specific features while learning new tasks. We validate REP’s superior resource efficiency over current state-of-the-art ViT- and CNN-based methods through extensive experiments on three image classification datasets.
@@ 1 Introduction
In this paper, we present several new techniques for reducing the computational and memory costs of prompt-based rehearsal-free methods that leverage frozen, pre-trained vision transformer (ViT) models [35, 36, 34, 29, 15]. Prompts are a small set of parameters that progressively learn incoming tasks to combat forgetting. Updates to these compact prompts incur minimum data writes without significantly harming the lifespan of device storage, which typically sustains up to 10K writes per location. This makes using prompts a good fit for on-device CL. Moreover, thanks to recent advancements in small-memory ViT models, such as ViT-Ti [30] (5.8M parameters; similar size to ResNet-10), we can always spot powerful ViT models that match the resource efficiency of traditional CNN models under consideration across memory specifications of typical resource-constrained IoT and embedded devices, as shown in Table 1.
@@
To enhance prompt-based CL for resource efficiency, we introduce REP (Resource-Efficient Prompting; Figure 1). REP is built upon our analysis of cost-accuracy trade-offs across the end-to-end learning process, with two key design insights. (1) The prompt selection stage, which constructs a prompt subset to augment input or intermediate data, is highly amenable to numerous promising options with fast approximations. (2) In contrast, the prompt update stage, which involves forward-backward passes over the backbone model, presents a range of optimizations with vastly different cost-accuracy trade-offs. With these insights, we develop three complementary techniques for both stages that trade minimal accuracy for high resource efficiency: (1) model downsizing for prompt selection and (2) adaptive token merging (AToM) and adaptive layer dropping (ALD) for prompt update. In pre-trained ViT models, task-specific knowledge is concentrated in shallow layers, while generalized feature information is spread across all layers [35] (See Figure 2 for our analysis). So, AToM and ALD carry out selective skipping across the data and model-layer dimensions in a “non-uniform” manner to preserve critical task-specific features in the shallow layers. REP provides resource efficiency that aligns with the capabilities of real-world commodity devices including NVIDIA Jetson TX2 and Nano. We evaluate REP across various memory budgets, backbone models, and datasets. Our experiments demonstrate that REP reduces wall-clock training time by 16–47% and lowers memory consumption by 5–33%.
@@ 1-4 GB L2P (ViT-B), DualP (ViT-B) MEMO (RN10, RN18)
Class-Incremental Learning. Our work primarily targets class-incremental learning (CIL) among various CL scenarios. In CIL, each incoming task introduces new classes within the same domain during the training phase, while task ID is not at hand during the inference phase [8]. In prior art using CNN backbone networks, rehearsal-based methods, which archive and replay representative samples from previous tasks [2, 4, 5, 28], have been a popular choice due to their superior performance [25]. Several recent methods are noteworthy along this line. For instance, BudgetCL [26] and CarM [19] suggest data-driven methods that outperform other strategies based on algorithmic optimizations, particularly when training is constrained by compute budgets. MEMO [40] effectively trades memory used to store old samples for saving task-specific model layers. We compare prompt-based methods with some of these methods and confirm better performance under resource constraints.
Continual Learning for the Edge. There has been a notable emphasis on efficient memory and energy usage in on-device learning, particularly in non-CL settings [9, 32, 33]. More recently, a handful of studies have explored extending CL capabilities to edge devices. Hayes and Kanan investigated online CL methods using CNN models tailored for embedded devices and provided valuable algorithmic insights. Similarly, Kwon et al. undertook a comparative analysis of rehearsal vs. regularization methods to uncover the cost-accuracy trade-offs, including storage, compute, and memory requirements. Ma et al. proposed an ML platform called Miro, which dynamically configures key design parameters of CNN-based CL methods to reduce energy costs within the device’s memory capacity. Unfortunately, none of these studies address the specific challenges associated with enabling vision transformers for on-device CL, which stands as a main goal of our research.
Prompting for Continual Learning. The use of prompts, which are explicit instructions or queries given to the model during training or inference, has shown promise in guiding continual learning for vision transformers [35, 36, 34, 29, 15, 13]. The general concept is to fine-tune prompts for new tasks while retaining knowledge acquired from previous tasks. L2P [36] and DualPrompt [35] propose prompt-tuning and prefix-tuning approaches, respectively, by selecting applicable prompts from a shared prompt pool. CODA-Prompt [29] introduces attention-conditioned prompts, which inherently facilitate higher prompting capacity. OVOR [13] uses only a single prompt to expedite the prompt selection stage. Our work suggests three complementary prompting techniques that can benefit all the above methods through full or even partial integration.
@@ 3.2.1 Unveiling System-Efficiency Opportunities
To answer this question, we explore the feature representations of pre-trained transformer blocks with prompts by analyzing the mean attention distance. Mean attention distance reflects the degree to which an attention head focuses on a specific patch relative to others in the input. This metric is widely used to analyze the layer representation of ViT models during the pre-training process [27, 7]. In our study, we re-purpose this concept to examine the behavior of frozen layers during data drift. For a given head in an attention layer within frozen blocks, let $(x_{0},y_{0})$ denote the position of a query patch, and $(x_{i},y_{i})$ denote the positions of patches it attends to, with attention weights $a_{i}$. The distance $d_{i}$—typically defined as the Euclidean or pixel distance—is calculated as $d_{i}=(x_{i}-x_{0})^{2}+(y_{i}-y_{0})^{2}$. We then calculate the weighted mean distance as $\frac{\sum_{i}a_{i}\cdot d_{i}}{\sum_{i}a_{i}}$.
@@ 3.2.2 Adaptive Token Merging (AToM)
Figure 3: The gradient norm for the prompt when applying conventional token merging (ToMe) vs AToM. L2P with ViT-L runs on Split ImageNet-R.
@@ 3.2.3 Adaptive Layer Dropping (ALD)
Inspired by insights from Figure 2 and prior work on progressive layer dropping (PLD) [38], we propose adaptive layer dropping (ALD; Algorithm 2) with two key features: (1) the dropping schedule considers both temporal and spatial dimensions [^1] The spatial dimension refers to the different layers of the model. Each layer processes input features at varying levels of abstraction., and (2) it manages to drop layers non-uniformly to preserve critical task-specific features in shallower layers. On the contrary, PLD only considers the temporal aspect and does not differentiate between layers when dropping. This results in poorer performance compared to ALD, as shown in Table 3.

### v3

@@ Abstract
Recent rehearsal-free methods, guided by prompts, excel in vision-related continual learning (CL) with drifting data but lack resource efficiency, making real-world deployment challenging. In this paper, we introduce Resource-Efficient Prompting (REP), which improves the computational and memory efficiency of prompt-based rehearsal-free methods while minimizing accuracy trade-offs. Our approach employs swift prompt selection to refine input data using a carefully provisioned model and introduces adaptive token merging (AToM) and layer dropping (ALD) for efficient prompt updates. AToM and ALD selectively skip data and model layers while preserving task-specific features during new task learning. Extensive experiments on multiple image classification datasets demonstrate REP’s superior resource efficiency over state-of-the-art ViT- and CNN-based methods.
@@ 1 Introduction
In this paper, we present several new techniques to reduce the computational and memory costs of prompt-based rehearsal-free methods that leverage frozen, pre-trained vision transformer (ViT) models [38, 39, 37, 31, 16]. Prompts are a small set of parameters that progressively learn incoming tasks to combat forgetting. Updates to these compact prompts incur minimum data writes without significantly harming the lifespan of device storage, which typically sustains up to 10K writes per location. This makes using prompts a good fit for on-device CL. Moreover, thanks to recent advancements in small-memory ViT models, such as ViT-Ti [32] (5.8M parameters; similar size to ResNet-10), we can always spot powerful ViT models that match the resource efficiency of traditional CNN models under consideration across memory specifications of typical resource-constrained IoT and embedded devices, as shown in Table 1.
@@
To enhance prompt-based CL for resource efficiency, we introduce REP (Resource-Efficient Prompting; Figure 1). REP is built upon our analysis of cost-accuracy trade-offs across the end-to-end learning process, with two key design insights. (1) The prompt selection stage, which constructs a prompt subset to augment input or intermediate data, is highly amenable to numerous promising options with fast approximations. (2) In contrast, the prompt update stage, which involves forward-backward passes over the backbone model, presents a range of optimizations with vastly different cost-accuracy trade-offs. With these insights, we develop three complementary techniques for both stages that trade minimal accuracy drop for high resource efficiency: (1) model downsizing for prompt selection and (2) adaptive token merging (AToM) and adaptive layer dropping (ALD) for prompt update. In pre-trained ViT models, task-specific knowledge is concentrated in shallow layers, while generalized feature information is spread across all layers [38] (See Figure 2 for our analysis). So, AToM and ALD carry out selective skipping across the data and model-layer dimensions in a “non-uniform” manner to preserve critical task-specific features in the shallow layers. REP provides resource efficiency that aligns with the capabilities of real-world commodity devices including NVIDIA Jetson TX2 and Nano. We evaluate REP across various memory budgets, backbone models, and datasets. Our experiments demonstrate that REP reduces training iteration time by up to 52% and lowers memory consumption by up to 32%.
@@ 1-4 GB L2P (ViT-B), DualP (ViT-B) MEMO (RN10, RN18)
Class-Incremental Learning: Our work primarily targets class-incremental learning (CIL) among various CL scenarios. In CIL, each incoming task introduces new classes within the same domain during the training phase, while task ID is not at hand during the inference phase [8]. In prior art using CNN backbone networks, rehearsal-based methods, which archive and replay representative samples from previous tasks [2, 4, 5, 30], have been a popular choice due to their superior performance [27]. Several recent methods are noteworthy along this line. For instance, BudgetCL [28] and CarM [21] suggest data-driven methods that outperform other strategies based on algorithmic optimizations, particularly when training is constrained by compute budgets. MEMO [43] effectively trades memory used to store old samples for saving task-specific model layers. We compare prompt-based methods with some of these methods and confirm better performance under resource constraints.
Continual Learning for the Edge: There has been a notable emphasis on efficient memory and energy usage in on-device learning, particularly in non-CL settings [9, 35, 36]. More recently, a handful of studies have explored extending CL capabilities to edge devices. Hayes and Kanan investigated online CL methods using CNN models tailored for embedded devices and provided valuable algorithmic insights. Similarly, Kwon et al. undertook a comparative analysis of rehearsal vs. regularization methods to uncover cost-accuracy trade-offs, including storage, compute, and memory requirements. Ma et al. proposed Miro, an ML platform that dynamically configures key design parameters of CNN-based CL methods to reduce energy costs within the device’s memory capacity. However, none of these studies address the challenges associated with enabling vision transformers for on-device CL, which stands as a main goal of our research.
Prompting for Continual Learning: The use of prompts, which are explicit instructions or queries given to the model during training or inference, has shown promise in guiding continual learning for vision transformers [38, 39, 37, 31, 16, 14]. The general concept is to fine-tune prompts for new tasks while retaining knowledge acquired from previous tasks. L2P [39] and DualPrompt [38] propose prompt-tuning and prefix-tuning approaches, respectively, by selecting applicable prompts from a shared prompt pool. CODA-Prompt [31] introduces attention-conditioned prompts, which inherently facilitate higher prompting capacity. OVOR [14] uses only a single prompt to expedite the prompt selection stage. Our work suggests three complementary prompting techniques that can benefit all the above methods through full or even partial integration.
@@ 3.2.1 Unveiling System-Efficiency Opportunities
To answer this question, we explore the feature representations of pre-trained transformer blocks with prompts by analyzing the mean attention distance. Mean attention distance reflects the degree to which an attention head focuses on a specific patch relative to others in the input. This metric is widely used to analyze the layer representation of ViT models during the pre-training process [29, 7]. In our study, we re-purpose this concept to examine the behavior of frozen layers during data drift.
For a given head in an attention layer within frozen blocks, let $(x_{0},y_{0})$ denote the position of a query patch, and $(x_{i},y_{i})$ the positions of patches that the attention head attends to, with corresponding attention weights $a_{i}$. The distance $d_{i}$—typically defined as the Euclidean or pixel distance—is calculated as $d_{i}=(x_{i}-x_{0})^{2}+(y_{i}-y_{0})^{2}$. We then calculate the weighted mean distance as $\frac{\sum_{i}a_{i}\cdot d_{i}}{\sum_{i}a_{i}}$.
@@ 3.2.2 Adaptive Token Merging (AToM)
Figure 3: The norm of gradient with respect to the prompt while training Split ImageNet-R when AToM and ToMe (Conventional Token Merging) is applied. ViT-L with L2P is used as a backbone.
@@ 3.2.3 Adaptive Layer Dropping (ALD)
Inspired by insights from Figure 2 and prior work on progressive layer dropping (PLD) [41], we propose adaptive layer dropping (ALD; Algorithm 2) with two key features: (1) the dropping schedule considers both temporal and spatial dimensions [^1] The spatial dimension refers to model layers, each processing input features at different levels of abstraction., and (2) it manages to drop layers non-uniformly to preserve critical task-specific features in shallower layers. On the contrary, PLD only considers the temporal aspect and does not differentiate between layers when dropping. This results in poorer performance compared to ALD, as shown in Table 4.
To strengthen our claim, we conduct a simple experiment that drops 25% of layers from various positions. We compare their training wall-clock GPU time and final average accuracy in Figure 4, using L2P with the ViT-L backbone on Split ImageNet-R. Specifically, Top-layer Drop and Bottom-layer Drop statically drop the first and last 25% of layers, respectively, while Random Drop randomly skips 25% of layers across the network. For comparison, we also include No Drop, which performs no layer dropping. Although all these naïve strategies incur accuracy loss, Top-layer Drop performs significantly worse than the others. Also, it appears undesirable to bypass only deep layers, even though these layers mainly capture general, task-agnostic features. While Random Drop most closely approaches No Drop’s accuracy, its uniform layer dropping still sacrifices some accuracy. This is why our method ALD carefully accounts for each layer’s spatial aspect in a non-uniform manner to minimize the impact on accuracy.
$\alpha(l)$ quantifies the degree of token merging performed by AToM in layer $l$. $n(l)$ represents the original number of tokens, and $n^{\prime}(l)$ is the number of tokens remaining after merging. When the number of merged tokens surpasses the threshold $\tau$, ALD adjusts the layer-dropping probability based on $\alpha$. Since deeper layers tend to merge more tokens with AToM, ALD is more likely to exceed $\tau$ in deeper layers and drop more aggressively. These parameters should be tuned to balance between efficiency and preserving the nuanced information contained in the merged tokens. We set $\alpha$ to 0.9, and $\tau$ to 16 for ViT-L, 12 for ViT-B, respectively.
One might contemplate exploiting stochastic depth [13], which gradually increases dropping probabilities for deeper layers without a temporal schedule. We compared ALD and stochastic depth and found that ALD generally offers 1.8–3.3% higher accuracy. This may be attributed to early training iterations after task insertion, which help stabilize training losses [23].
Resource Efficiency: REP triggers AToM and ALD for each task insertion to optimize both time and memory usage. Much like our prompt selection optimization, AToM consistently enhances time and memory efficiency. In contrast, ALD solely contributes to reducing time costs because it operates layer-keeping probability $\theta_{t,l}$ starting at 1.0, i.e., no layer dropping, which necessitates traversing full layers.
4 Experiments
Data Generation: To organize task streams, we use three image classification datasets: CIFAR-100 (100 classes) [18], ImageNet-R (200 classes) [12], and PlantDisease (38 classes) [26]. Out of these datasets, ImageNet-R is known for exhibiting much higher intra-class variability among images and an uneven class-size distribution. We divide CIFAR-100 and ImageNet-R into 10 tasks to create Split CIFAR-100 (i.e., 10 classes per task) and Split ImageNet-R (i.e., 20 classes per task), respectively [39, 38, 31]. For PlantDisease, we drop 3 plant disease classes with very few images and organize the remaining 35 classes into 7 tasks to create Split PlantDisease (i.e., 5 classes per task).
Methods: We employ L2P [39], DualPrompt [38], and CODA-Prompt [31] as representative ViT-based prompting methods. L2P and DualPrompt select suitable prompts via inference, similar to [34], while CODA-Prompt employs a strategy for adaptively optimizing prompts without prompt selection, similar to [14]. These methods capitalize on ImageNet pre-trained models as backbones: ViT-L (307M), ViT-B (86M), and ViT-Ti (5.8M)—ViT-Ti is one of the smallest vision transformer models to our knowledge. We categorize the prompt-based methods into these three backbone models.
To demonstrate the effectiveness of prompting methods, we also incorporate two state-of-the-art CNN-based methods, BudgetCL [6] and MEMO [43]. Both BudgetCL and MEMO improve model accuracy by leveraging spare memory to store and replay past samples, with MEMO additionally storing model checkpoints from history. These methods also use ImageNet pre-trained models as backbones: ResNet-10 (RN10; 5M), ResNet-18 (RN18; 11M), ResNet-26 (RN26; 14M), ResNet-34 (RN34; 22M), and ResNet-101 (RN101; 43M) models. For fair comparisons, we similarly group methods based on their memory usage. The analysis result is presented in  Section 4.1. See  Appendix B for detailed settings of all methods mentioned above.
Hardware and Metrics: We use NVIDIA RTX 3090 as a reference GPU to cover a wide range of memory sizes while maintaining consistent computational power and measure the GPU iteration time required for computation—GPU iteration time can be correlated linearly with energy usage [23]. In Appendix H, we evaluate our memory-efficient method directly on NVIDIA Jetson TX2, although it is applicable beyond this specific device. We also report the final average accuracy by averaging the accuracy of all classes after the last task training [5, 31], which assesses the decline in task accuracy averaged over all tasks.
4.1 Preliminary Empirical Study
Table 2: Accuracy and computational cost of all competing methods on Split CIFAR-100, Split ImageNet-R, and Split PlantDisease datasets. We report Final Average Accuracy (Acc.) both without and with REP. Iteration time (ms) and memory usage (GB) are provided in absolute terms, along with their percentage values relative to REP. ($\uparrow$) higher is better, ($\downarrow$) lower is better.
Model Method Split CIFAR-100 Split ImageNet-R Split PlantDisease Iter. time (ms) Mem. (GB)
Acc. ($\uparrow$) w/ REP Acc. ($\uparrow$) w/ REP Acc. ($\uparrow$) w/ REP Iter.time ($\downarrow$) w/ REP Mem. ($\downarrow$) w/ REP
ViT-L L2P 88.2$\pm$0.3 86.4$\pm$0.3 75.6$\pm$1.0 75.3$\pm$0.9 75.9$\pm$3.1 81.1$\pm$3.9 446.9 235.4 (-47.3%) 6.5 4.5 (-31.5%)
DualPrompt 86.3$\pm$0.3 84.5$\pm$0.6 71.2$\pm$0.6 70.6$\pm$0.9 75.1$\pm$1.1 78.2$\pm$2.9 423.7 202.5 (-52.2%) 5.9 4.3 (-27.8%)
CODA-Prompt 85.4$\pm$0.4 83.7$\pm$0.8 74.4$\pm$0.6 72.4$\pm$0.9 75.0$\pm$2.0 74.0$\pm$2.7 568.0 431.1 (-24.1%) 13.2 11.0 (-16.8%)
UpperBound 94.1$\pm$0.1 - 85.2$\pm$0.2 - 99.6$\pm$0.3 - 426.7 - 9.8 -
ViT-B L2P 84.5$\pm$0.8 83.4$\pm$1.8 59.4$\pm$0.8 59.7$\pm$1.2 64.4$\pm$2.6 64.0$\pm$2.4 142.5 100.1 (-29.8%) 2.3 2.0 (-11.8%)
DualPrompt 85.1$\pm$0.2 84.9$\pm$0.9 68.1$\pm$0.1 67.0$\pm$1.0 78.0$\pm$1.5 77.0$\pm$1.8 133.4 99.7 (-25.2%) 2.1 1.8 (-11.2%)
CODA-Prompt 83.3$\pm$0.9 81.3$\pm$1.9 59.0$\pm$0.3 58.0$\pm$1.7 71.4$\pm$0.8 70.2$\pm$2.7 164.1 144.8 (-11.8%) 6.9 5.7 (-17.8%)
UpperBound 92.0$\pm$0.2 - 81.4$\pm$0.1 - 99.6$\pm$0.4 - 143.2 - 3.2 -
ViT-Ti L2P 60.3$\pm$0.2 58.9$\pm$1.3 41.2$\pm$0.7 40.3$\pm$1.0 56.1$\pm$2.9 55.0$\pm$2.8 33.8 29.0 (-14.2%) 0.5 0.5 (-2.1%)
DualPrompt 62.9$\pm$0.5 62.5$\pm$0.9 43.6$\pm$0.8 42.7$\pm$1.0 64.0$\pm$1.0 61.9$\pm$1.8 33.9 30.0 (-11.6%) 0.4 0.4 (-4.8%)
CODA-Prompt 67.0$\pm$0.4 65.9$\pm$1.8 49.3$\pm$0.8 47.6$\pm$2.0 66.8$\pm$0.8 65.0$\pm$2.6 36.1 29.8 (-17.4%) 2.8 2.8 (-0.7%)
UpperBound 83.0$\pm$0.1 - 77.4$\pm$0.2 - 98.0$\pm$0.9 - 42.3 - 0.6 -
Scaling Across Varying Backbone Networks.  ViT-based methods scale well, with larger backbone networks overall yielding higher accuracy. In contrast, CNN-based methods scale poorly with increased backbone size. Specifically, on Split ImageNet-R, a more challenging dataset, MEMO/RN34 is only 2.95% better than MEMO/RN18, despite being around twice larger. On Split CIFAR-100, increasing the backbone size yields only limited gains.
Memory Efficiency.  CNN-based methods are often considered more memory-efficient because they use backbones a magnitude smaller than ViT models. Thus, it is commonly believed that ample memory can be allocated to replay samples or past model checkpoints [43], which help improve model performance. However, in CL, which involves typical training iterations, a substantial amount of memory must be allocated for feature maps [22], leaving little room for memory buffers. For instance, for a device memory range of 1–4GB in Table 1, the backbone of MEMO/RN18 consumes only 43MB of memory (6.5% of ViT-L), but feature maps occupy as much as 1.7GB (more than ViT-based methods). This causes MEMO/RN18 to be short on memory for the replay buffer, dramatically degrading model accuracy, as observed in Figure 5.
4.2 Main Results
Table 2 presents the efficacy of REP when integrated into prompt-based CL methods across various ViT backbones (ViT-L, ViT-B, ViT-Ti) and datasets.
Resource Efficiency:  Integrating REP into existing prompt-based CL methods significantly reduces computational overhead and memory usage across all model configurations: we observe memory savings of up to 32% for ViT-L, 18% for ViT-B, and 5% for ViT-Ti.
Although the backbone network parameters remain frozen during training, the model must still perform a backward pass to compute losses for both classification and prompt learning. As a result, a substantial amount of memory is allocated to store intermediate activations. By selectively merging tokens (AToM) and adaptively dropping layers (ALD), REP mitigates these memory demands without compromising performance-critical information flow.
The advantages of REP extend to training speed, which is closely linked to energy consumption. For ViT-L, the largest architecture considered, REP reduces iteration time by up to 52% compared to baseline methods, translating directly into lower energy consumption. In contrast, with the more compact ViT-Ti backbone, resource savings persist but are less pronounced due to the model’s already low computational cost. These results highlight the versatility of REP, demonstrating its scalability from compact to large models while ensuring substantial resource efficiency gains across different regimes.
Accuracy:  While L2P, DualPrompt, and CODA-Prompt were originally designed and evaluated on a ViT-B backbone, our study systematically extends these methods across a diverse range of ViT architectures, considering their energy-accuracy trade-offs in on-device CL under various memory constraints. This broader perspective reveals notable gains in resource efficiency with REP, archived with only marginal accuracy drops, e.g., 0.03–1.36% on Split CIFAR-100 and 0.26–1.62% on Split ImageNet-R. In some cases, accuracy even improves. For instance, on ViT-L with Split PlantDisease, L2P’s accuracy increases from 75.92% to 81.08% when REP is applied. These findings suggest that AToM and ALD effectively eliminate redundant computations while preserving essential features for CL.
Table 3: Ablation study of REP components, demonstrating their contributions to both resource efficiency and accuracy.
REP (ours) 75.6$\pm$1.0 3.6$\pm$0.3 2542.4 4.5
(1) $f_{\text{efficient}}$ 74.8$\pm$0.8 4.0$\pm$0.8 3698.8 5.5
(2) AToM 74.9$\pm$0.8 3.5$\pm$0.8 4442.0 5.6
(3) ALD 74.5$\pm$0.8 3.6$\pm$0.7 4244.0 6.7
(4) $f_{\text{efficient}}$ + AToM 74.2$\pm$0.8 4.0$\pm$0.8 2861.4 4.5
(5) $f_{\text{efficient}}$ + ALD 74.5$\pm$0.4 3.4$\pm$0.5 3206.7 5.5
(6) AToM + ALD 74.6$\pm$0.4 2.6$\pm$0.5 3448.8 4.8
4.3 Ablation and Additional Study
Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
REP (ours) 75.6$\pm$1.0 3.6$\pm$0.3 2542.4 4.5
(1) w/ ToMe 70.2$\pm$0.7 2.6$\pm$0.9 2917.8 3.7
(2) w/ PLD 73.3$\pm$0.7 3.9$\pm$0.7 2741.9 4.5
2 76.2$\pm$0.7 3.5$\pm$0.7 2808.4 5.5
4 75.3$\pm$0.3 3.3$\pm$0.5 2714.0 5.2
8 75.6$\pm$0.9 3.6$\pm$0.3 2542.4 4.5
0.25 73.2$\pm$0.6 4.0$\pm$0.5 2362.1 4.5
0.50 75.6$\pm$1.0 3.6$\pm$0.3 2542.4 4.5
0.75 74.4$\pm$0.6 3.5$\pm$0.5 2861.4 4.5
Component Ablation: In  Table 3, we ablate REP ’s components to assess their contributions to resource efficiency. Each component plays a key role in reducing memory and compute time, with ALD affecting compute time only, as expected. Notably, AToM substantially optimizes both resource usages.
Algorithm Validation: Token merging (ToMe) [3] and progressive layer dropping (PLD) [41] are specifically designed to accelerate traditional transformer-based model training. To validate the importance of incorporating our adaptive techniques instead in CL, we evaluate how REP performs in case AToM and ALD are replaced with ToMe and PLD, respectively. The results are presented in Table 4. Although applying ToMe or PLD improves system efficiency over L2P, it results in an accuracy loss of 5.15% or 2.01%, respectively, compared to using our techniques. This indicates that ToMe and PLD are less desirable for achieving our goal, i.e., enhancing system efficiency without compromising accuracy.
AToM and ALD Intensity: Table 5 shows the effects of varying intensities of AToM and ALD, focusing on the number of merged tokens ($n$) in AToM and the keep ratio ($\theta$) in ALD. AToM appears to maintain stable accuracy even as more tokens are merged. In contrast, for ALD, a lower keep ratio improves system efficiency by reducing iteration time and memory usage, but it can markedly impair model accuracy if the ratio is too low. Overall, when used with caution, AToM and ALD can effectively balance system efficiency and accuracy.
5 Conclusion
In this work, we introduced Resource-Efficient Prompting (REP), a framework designed to enhance the computational and memory efficiency of prompt-based rehearsal-free continual learning methods. By incorporating adaptive token merging and adaptive layer dropping, REP selectively streamlines the learning process while preserving task-specific features, significantly reducing resource consumption without compromising accuracy. Experiments on multiple image classification datasets demonstrate that REP achieves substantial efficiency gains over state-of-the-art ViT- and CNN-based methods, making it a practical solution for on-device continual learning.
Since our study focuses on scenarios wherein a sequence of tasks is presented within learning environments, our method does not determine the curriculum of learning. Future work could extend our our method for curriculum-based continual learning. Moreover, although using replay buffer may not be feasible in real-world scenarios due to data privacy concerns, on-device CL is already viewed as a privacy-preserving approach, making rehearsal-based REP a promising direction.
Appendix A Algorithm Details